Database last updated on Mon, 02 Feb 2026 00:40:50 +0000 by localhost
Instrument constant changes for the week before 2023299 (1 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 20 rows)
Measurements and tests summary.
| _jday | _loc | _serial | _skc | aode | ds | dt | ed | fm | hg | rs | sl | uv | zs |
| 299 | TORONTO | #014 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.622) | 3 | 3 | 1 | 2 | 42 | 16 | 1 | 6 | 48 | 64 |
(1 sec)
Number of days since last dome/window cleaning (0 sec, 2 rows)
| serial | days |
| 014 | more than a year |
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 014 | 2023300 | 0-1 |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 014 | 14 | 1 | 6 | 1 | 1 | 1 | 1 | 1 | 1 |
AP results for MB Brewer. (0 sec, 1 rows)
| Serial | Jday | t1 off | t1 on | t2 off | t2 on | t3 off | t3 on | HV off | HV on | +15V off | +15V on | +5V off | +5V on | -15V off | -15V on | +24V off | +24V on | +5V sps off | +5V sps on | -8V sps off | -8V sps on | SL A off | SL A on | SL V off | SL V on | W |
| 014 | 2023299 | 30.6 | 30.5 | 26.2 | 26.2 | 29.3 | 29.3 | 1554.3 | 1548.0 | 15.1 | 15.0 | 5.1 | 5.1 | -15.1 | -15.1 | 24.1 | 24.1 | 4.9 | 5.0 | -7.9 | -7.9 | 0.1 | 1.6 | 0.0 | 10.93 | 17.3 |
Standard lamp intensity analysis (0 sec, 1 rows)
| Serial | min | max | filter | fixable | 30-day history |
| #014 | 2962330 | 2962330 | 0 | yes |